# From Cradle to Cane: A Two-Pass Framework for High-Fidelity Lifespan Face Aging

### v1

@@ Abstract
Face aging has become a crucial task in computer vision, with applications ranging from entertainment to healthcare. However, existing methods struggle with achieving a realistic and seamless transformation across the entire lifespan, especially when handling large age gaps or extreme head poses. The core challenge lies in balancing age accuracy and identity preservation—what we refer to as the Age-ID trade-off. Most prior methods either prioritize age transformation at the expense of identity consistency or vice versa. In this work, we address this issue by proposing a two-pass face aging framework, named Cradle2Cane, based on few-step text-to-image (T2I) diffusion models. The first pass focuses on solving age accuracy by introducing an adaptive noise injection (AdaNI) mechanism. This mechanism is guided by including prompt descriptions of age and gender for the given person as the textual condition. Also, by adjusting the noise level, we can control the strength of aging while allowing more flexibility in transforming the face. However, identity preservation is weakly ensured here to facilitate stronger age transformations. In the second pass, we enhance identity preservation while maintaining age-specific features by conditioning the model on two identity-aware embeddings (IDEmb): SVR-ArcFace and Rotate-CLIP. This pass allows for denoising the transformed image from the first pass, ensuring stronger identity preservation without compromising the aging accuracy. Both passes are jointly trained in an end-to-end way. Extensive experiments on the CelebA-HQ test dataset, evaluated through Face++ and Qwen-VL protocols, show that our Cradle2Cane outperforms existing face aging methods in age accuracy and identity consistency. Additionally, Cradle2Cane demonstrates superior robustness when applied to in-the-wild human face images, where prior methods often fail. This significantly broadens its applicability to more diverse and unconstrained real-world scenarios.
@@ 1 Introduction
Deep learning lecun2015deeplearning has allowed a realistic alteration of the apparent age of a person chen2023face ; deng2019arcface ; zhang2017age , opening promising applications in areas such as computer graphics, entertainment, forensics and healthcare. The goal of facial age transformation is to simulate the natural aging or de-aging process in a visually convincing manner. To this end, numerous methods have been developed to achieve high-quality, identity-preserving age progression and regression. Recent approaches are based on deep generative models, such as generative adversarial networks (GANs) alaluf2021only ; goodfellow2014gan ; hsu2022agetransgan and Diffusion Models (DMs) banerjee2023identity ; chen2023face ; ito2025selfage ; wahid2024diffage3d , and have shown promising results. However, to the best of our knowledge, most existing methods suffer from a limited transformation range and often struggle to maintain high-fidelity results when handling large age gaps, occlusions, or extreme head poses. As a result, they fall short of delivering seamless cradle-to-cane face aging performance across the entire lifespan.
@@
In this study, we attribute the limitations of existing face aging methods to an imbalanced trade-off between age accuracy and identity consistency—a challenge we term the Age-ID trade-off. Most prior approaches alaluf2021only ; chen2023face ; ito2025selfage tend to emphasize one aspect while neglecting the other due to their unified framework to deal with entire lifespan ages, resulting in either visually convincing age transformations that compromise identity, or identity-preserving outputs with inaccurate aging effects. This imbalance is evident in the trade-off curves shown in Fig. 1, where, for example, as the age difference increases, methods such as FADING chen2023face , CUSP gomez2022custom , and Lifespan or2020lifespan tend to show improved aging realism at the cost of reduced identity preservation, or vice versa. The fluctuating curves of harmonic means further demonstrate this phenomenon. To address the Age-ID trade-off problem, we propose our face aging framework built upon few-step T2I diffusion models SDXL-Turbo sauer2024adversarial , which offer two key advantages: 1) the few-step nature of these models dao2024swiftbrushv2 ; li2023faster ; liu2023instaflow ; sauer2024adversarial enables fast inference while maintaining high image fidelity, and 2) the flexible noise control in the forward diffusion process allows fine-grained modulation of aging strength by adjusting the injected noise scale. As illustrated in Fig. 2, injecting higher noise levels during the forward diffusion process increases editability, enabling more pronounced aging transformations while downgrading the identity consistency. In contrast, lower noise levels better preserve identity information with less aging accuracy, highlighting the trade-off between visual age change and identity consistency. Similar identity-editing trade-offs are also observed in image editing and generation hertz2022prompt ; li2023stylediffusion ; meng2021sdedit ; kai2023DPL . However, directly applying the few-step diffusion models cannot achieve fine-grained face aging with identity consistency and age accuracy, resulting from that the few-step T2I diffusion models ho2022imagen ; Rombach_2022_CVPR_stablediffusion ; deepfloyd do not inherently support age or identity conditions.
@@ 4.2 Experimental Results
Method Type Face++ Evaluation (CelebA-HQ) Qwen-VL Evaluation (CelebA-HQ) Qwen-VL (CelebA-HQ-in-the-wild) Inference Time (s) Train Data
@@ 5 Conclusion
Appendix
Appendix A Limitations
Appendix B Broader Impacts
Appendix C Experiments details
C.1 Implementation Details.
[…]
C.2 Prompts for Qwen-VL Evaluation
Age Estimation
Image Quality
Identity Similarity
C.3 Harmonic Consistency Score (HCS)
C.4 Open-Source Implementations and Settings
Method Repository Link
SAM [1] https://github.com/yuval-alaluf/SAM
FADING [6] https://github.com/MunchkinChen/FADING
CUSP [13] https://github.com/guillermogotre/CUSP
HRFAE [69] https://github.com/InterDigitalInc/HRFAE
C.5 Age–ID Trade-off Evaluation Details
[…]
C.6 User Study Details
Appendix D Additional Results
D.1 Qualitative Comparison on in-the-wild Images
D.2 Generalization to Diverse Reference Ages
D.3 Face Aging across the Entire Lifespan

### v2

@@ Abstract
Face aging has become a crucial task in computer vision, with applications ranging from entertainment to healthcare. However, existing methods struggle with achieving a realistic and seamless transformation across the entire lifespan, especially when handling large age gaps or extreme head poses. The core challenge lies in balancing age accuracy and identity preservation—what we refer to as the Age-ID trade-off. Most prior methods either prioritize age transformation at the expense of identity consistency or vice versa. In this work, we address this issue by proposing a two-pass face aging framework, named Cradle2Cane, based on few-step text-to-image (T2I) diffusion models. The first pass focuses on solving age accuracy by introducing an adaptive noise injection (AdaNI) mechanism. This mechanism is guided by including prompt descriptions of age and gender for the given person as the textual condition. Also, by adjusting the noise level, we can control the strength of aging while allowing more flexibility in transforming the face. However, identity preservation is weakly ensured here to facilitate stronger age transformations. In the second pass, we enhance identity preservation while maintaining age-specific features by conditioning the model on two identity-aware embeddings (IDEmb): SVR-ArcFace and Rotate-CLIP. This pass allows for denoising the transformed image from the first pass, ensuring stronger identity preservation without compromising the aging accuracy. Both passes are jointly trained in an end-to-end way. Extensive experiments on the CelebA-HQ test dataset, evaluated through Face++ and Qwen-VL protocols, show that our Cradle2Cane outperforms existing face aging methods in age accuracy and identity consistency. Additionally, Cradle2Cane demonstrates superior robustness when applied to in-the-wild human face images, where prior methods often fail. This significantly broadens its applicability to more diverse and unconstrained real-world scenarios. Code is available at https://github.com/byliutao/Cradle2Cane.
@@ 1 Introduction
Deep learning lecun2015deeplearning has allowed a realistic alteration of the apparent age of a person chen2023face ; deng2019arcface ; zhang2017age , opening promising applications in areas such as computer graphics, entertainment, forensics and healthcare. The goal of facial age transformation is to simulate the natural aging or de-aging process in a visually convincing manner. To this end, numerous methods have been developed to achieve high-quality, identity-preserving age progression and regression. Recent approaches are based on deep generative models, such as generative adversarial networks (GANs) alaluf2021only ; goodfellow2014gan ; hsu2022agetransgan and Diffusion Models (DMs) banerjee2023identity ; chen2023face ; ito2025selfage ; wahid2024diffage3d ; wu2025representationentanglementgenerationtraining , and have shown promising results. However, to the best of our knowledge, most existing methods suffer from a limited transformation range and often struggle to maintain high-fidelity results when handling large age gaps, occlusions, or extreme head poses. As a result, they fall short of delivering seamless cradle-to-cane face aging performance across the entire lifespan.
@@
In this study, we attribute the limitations of existing face aging methods to an imbalanced trade-off between age accuracy and identity consistency—a challenge we term the Age-ID trade-off. Most prior approaches alaluf2021only ; chen2023face ; ito2025selfage tend to emphasize one aspect while neglecting the other due to their unified framework to deal with entire lifespan ages, resulting in either visually convincing age transformations that compromise identity, or identity-preserving outputs with inaccurate aging effects. This imbalance is evident in the trade-off curves shown in Fig. 1, where, for example, as the age difference increases, methods such as FADING chen2023face , CUSP gomez2022custom , and Lifespan or2020lifespan tend to show improved aging realism at the cost of reduced identity preservation, or vice versa. The fluctuating curves of harmonic means further demonstrate this phenomenon. To address the Age-ID trade-off problem, we propose our face aging framework built upon few-step T2I diffusion models SDXL-Turbo sauer2024adversarial , which offer two key advantages: 1) the few-step nature of these models dao2024swiftbrushv2 ; li2023faster ; liu2023instaflow ; sauer2024adversarial enables fast inference while maintaining high image fidelity, and 2) the flexible noise control in the forward diffusion process allows fine-grained modulation of aging strength by adjusting the injected noise scale. As illustrated in Fig. 2, injecting higher noise levels during the forward diffusion process increases editability, enabling more pronounced aging transformations while downgrading the identity consistency. In contrast, lower noise levels better preserve identity information with less aging accuracy, highlighting the trade-off between visual age change and identity consistency. Similar identity-editing trade-offs are also observed in image editing and generation hertz2022prompt ; li2023stylediffusion ; meng2021sdedit ; kai2023DPL ; tang2024locinv ; hu2025anchor_token . However, directly applying the few-step diffusion models cannot achieve fine-grained face aging with identity consistency and age accuracy, resulting from that the few-step T2I diffusion models ho2022imagen ; Rombach_2022_CVPR_stablediffusion ; deepfloyd ; li2025oneway_ticket do not inherently support age or identity conditions.
@@ 4.2 Experimental Results
Method Type Face++ Evaluation (CelebA-HQ) Qwen-VL Evaluation (CelebA-HQ) Qwen-VL (CelebA-HQ-in-the-wild)
Inference
Time (s)
Train
Data
@@ 5 Conclusion
Acknowledgements
This research was supported by the collaborative project between Beijing Samsung Telecommunications Technology Co., Ltd. and the Tianjin Key Laboratory of Visual Computing and Intelligent Perception (VCIP) of Nankai University, entitled “Generated Face for Enhancing Face Dataset” (Project No. SRC-Beijing-DVL-2024-00241).
We would like to express our sincere gratitude to all co-authors for their invaluable contributions and insightful suggestions. We are particularly grateful to Yaxing Wang (Associate Professor, Nankai University) and Kai Wang (Assistant Professor, City University of Hong Kong (Dongguan)). Their meticulous advice and guidance were instrumental in the completion of this research.
@@ 80. Guidelines:
Appendix
Appendix A Limitations
Appendix B Broader Impacts
Appendix C Experiments details
C.1 Implementation Details.
Our framework is built upon the SDXL-Turbo architecture and trained on the FFHQ dataset, which contains high-quality facial images annotated with age and gender labels. To improve background consistency, we employ CarveKit [60] for foreground-background segmentation, replacing non-facial regions with a uniform gray mask. During training, we guide the model using text prompts in the format: "a face image of a {target age} years old {gender}" where gender is either female or male. All images are generated at a fixed resolution of $\times512$.
C.2 Prompts for Qwen-VL Evaluation
Age Estimation
Image Quality
Identity Similarity
C.3 Harmonic Consistency Score (HCS)
C.4 Open-Source Implementations and Settings
Method Repository Link
SAM [1] https://github.com/yuval-alaluf/SAM
FADING [7] https://github.com/MunchkinChen/FADING
CUSP [14] https://github.com/guillermogotre/CUSP
HRFAE [76] https://github.com/InterDigitalInc/HRFAE
C.5 Age–ID Trade-off Evaluation Details
To quantitatively evaluate the trade-off between age accuracy and identity consistency, we selected 50 male and 50 female face images from the CelebA-HQ test dataset. For each image, we generated aging results across age offsets ranging from 60-60 to +60+60 with a step size of $1$ year, excluding the zero offset. Identity similarity was measured using ArcFace by computing the cosine similarity between the original and age-edited images. Age similarity was quantified based on the predicted age error from the MiVOLO estimator. Specifically, we computed the age consistency score as $\text{age\_cosine}=1-\frac{|\text{age}_{\text{pred}}-\text{age}_{\text{target}}|}{\text{max\_age\_diff}}$, where max_age_diff is set to $40$. This score ranges from 0 to $1$, with higher values indicating better age alignment. Comparisons with the remaining methods are illustrated in Fig. 7. Note that IPFE [3] does not support continuous year-level age control, thus only the provided age groups reported in the original paper were included in our evaluation.
C.6 User Study Details
Appendix D Additional Results
D.1 Qualitative Comparison on in-the-wild Images
D.2 Generalization to Diverse Reference Ages
D.3 Face Aging across the Entire Lifespan
Appendix E Additional Experiment
E.1 Additional Quantitative Experiment
To further validate the effectiveness of our approach, we conducted additional quantitative experiments on extra datasets and baselines. Specifically, we compared our method with recent generic face-editing systems (e.g., StyleCLIP [50] and FaceDNeRF [78]) on the standard aging datasets AgeDB-30 [46], CACD [6], and FG-NET. For each dataset, we generated approximately 200 images per method and evaluated them using four metrics: age difference (Age Diff.), identity similarity (ID Sim.), image quality (Img. Quality), and Holistic Consistency Scores (HCS). The comprehensive results are summarized in Table 4, where lower Age Diff. and higher scores on the remaining metrics indicate better performance.
Method AgeDB-30 CACD FG-NET
CUSP [14] 8.29 76.30 82.80 77.75 10.80 75.40 85.86 74.18 19.45 74.20 79.79 60.71
SAM [1] 8.72 58.91 86.57 67.19 9.25 67.27 87.57 71.75 5.33 72.18 81.98 78.77
FADING [7] 8.10 76.00 79.22 77.82 5.16 71.27 86.44 78.39 7.22 73.15 79.43 77.30
Styleclip [50] 14.73 62.12 86.60 62.64 16.28 70.64 87.80 64.48 14.73 73.82 88.19 68.08
Facednerf [78] 12.60 53.78 91.97 60.25 13.05 59.00 92.02 62.91 13.21 56.37 92.43 61.21
Cradle2Cane 5.57 73.00 90.01 79.00 4.63 66.73 90.02 76.06 5.79 67.18 85.03 75.25
Table 4: Comparison on AgeDB-30, CACD and FG-NET datasets. Best results are marked in blue, and second-best in green.
Our method, Cradle2Cane, demonstrates exceptional performance across all three benchmarks. On AgeDB-30, it achieves state-of-the-art results, securing the best age accuracy and the top harmonic score, which underscores its superior overall balance. This strong performance continues on the CACD dataset, where our method again delivers the best age accuracy while maintaining highly competitive scores on other metrics. Even on the highly challenging FG-NET dataset, Cradle2Cane proves its robustness by delivering the second-best age accuracy and demonstrating strong, consistent performance against all competitors.
E.2 Ablation Study on Threshold Selection
To investigate the rationale behind our choice of age thresholds, we conducted an ablation study with multiple division settings. Specifically, we compared thresholds of {5,15}, {5,20}, {7,22}, {10,30}, {15,35}, and {20,40}. The results presented in Table 5 reveal a distinct trade-off between age accuracy and identity preservation. We observe that wider threshold intervals, such as {15,35} and {20,40}, yield superior identity similarity (ID Sim.) and Holistic Consistency Scores (HCS). However, this gain is achieved at the expense of age fidelity, as evidenced by the significant increase in the Age Difference metric from a low of 4.92 to 6.71.
Given that age accuracy is the central objective of our work, we selected the {5, 20}division as our default configuration. This setting achieves the best performance in age accuracy while maintaining a competitive balance in identity preservation and image quality. This choice ensures our primary goal is met and establishes a rigorous, transparent baseline for our experiments.
Table 5: Ablation study on different threshold divisions. This analysis highlights the trade-off between age accuracy and identity preservation. Best results are marked in blue, and second-best in green.
Thresholds Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑
{5, 15} 4.93 73.18 92.28 79.77
{5, 20} 4.92 73.73 92.60 80.11
{7, 22} 5.06 74.82 92.89 80.60
{10, 30} 5.46 76.91 92.56 81.36
{15, 35} 5.64 77.09 92.82 81.26
{20, 40} 6.71 78.55 92.74 80.82
E.3 Ablation Study on Robustness and Architectural Contributions
To verify whether the robustness of our method on in-the-wild face images primarily comes from background removal pre-processing (Carvekit) or from the model architecture itself, we conducted an ablation study. Specifically, we compared our method with and without Carvekit pre-processing. Both variants were trained for 10 epochs on 1,000 images and evaluated on the CelebA-in-the-wild dataset. The results are shown in Table 6.
Table 6: Ablation study of the background removal pre-processing (Carvekit). The minor performance difference highlights that robustness is intrinsic to the model architecture. Best results are marked in blue.
Method Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑
w/ Carvekit 6.76 62.00 91.75 71.02
w/o Carvekit 7.00 62.27 91.60 70.97
As presented in the table, the performance impact of background removal is marginal. This finding indicates that Carvekit serves as a beneficial but non-essential preprocessing step, rather than the primary source of the model’s robustness. Instead, the method’s resilience to in-the-wild variations is primarily attributed to its architectural design. First, the SDXL-Turbo backbone, pre-trained on a vast and diverse dataset, provides a strong foundation for generalization across varied poses, lighting conditions, and expressions. Second, our proposed IDEmb module systematically reinforces identity preservation. The SVR-ArcFace component extracts a stable identity embedding that is disentangled from transient attributes, while Rotate-CLIP executes minimal, precise modifications within CLIP’s robust semantic space. This dual mechanism ensures that the age attribute is altered while other original characteristics are preserved with high fidelity. In summary, this study confirms that our model’s robustness is an intrinsic property of its architecture, with SDXL-Turbo enabling generalization and IDEmb ensuring identity-consistent editing.
Algorithm 1 The Proposed Cradle2Cane Framework
1:Input: Source image $\mathbf{x}_{a}$, source age $a$, target age $b$.
2:Output: Final aged image $\mathbf{x}_{b}$.
3:
4:// Pass 1: Adaptive Age Transformation
5:Select noise level $\mathbf{z}_{i}$ based on $|\Delta\text{age}|$ per (Eq. 1).
[…]
7:Denoise from noise level $\mathbf{z}_{i}$ with condition $c_{\text{age}}$ to get latent $\hat{\mathbf{z}}_{0}$.
[…]
9:
10:// Pass 2: Identity Enhancement
11:// — IDEmb Generation —
12:Generate aged variants $\{\mathbf{x}_{b}^{(i)}\}_{i=1}^{n}$ from $\mathbf{x}_{a}$.
[…]
19:
[…]
23:
24:// — Final Refinement —
[…]
27:Set low noise timestep $T_{\text{low}}$.
[…]
29:for $t=T_{\text{low}},\dots,1$ do $\mathbf{z}_{t-1}\leftarrow p_{\theta}(\mathbf{z}_{t},t,c_{\text{ID}})$ $\triangleright$ Denoise with identity condition
30:end for
[…]
32:
33:Return $\mathbf{x}_{b}$.